OpenStudy (anonymous):

1/4 l4x+7l = 8x + 16 Solve for this equation. Now the answer is -71/36 but, I don't know how they solved to get that answer. Can someone help?

OpenStudy (anonymous):

@monokerous any idea how to solve it?

OpenStudy (phi):

change the absolute value to parens and solve. also solve -(1/4) (4x+7) = 8x+16 check both solutions to make sure they are allowed.

OpenStudy (anonymous):

@phi check both solutions? what do you mean?

OpenStudy (anonymous):

For absolute values, there are always two solutions that you get.

OpenStudy (anonymous):

I know so how would i set up both equations to get my 2 solutions?

OpenStudy (anonymous):

I'm trying to look for a useful site (as a refresher or something) as I've only ever learnt absolute values once, briefly. lol. Sorry.

OpenStudy (anonymous):

I tried doing that but, i didn't have much luck it doesnt give an example like this

OpenStudy (anonymous):

Agreed.

OpenStudy (phi):

the first equation: if the stuff inside the abs value is positive, you can drop the | |: (1/4) ( 4x+7)= 8x+16 solve for x if the stuff inside is negative, then we can say (-1/4)(4x+7)= 8x+16 solve for x the solution may be invalid for one of the x. Try it.

OpenStudy (anonymous):

Ok, I will solve it. Thanks I will let you know what i get

OpenStudy (anonymous):

would i distribute (1/4) with (4x +7)??

OpenStudy (phi):

1/4 l4x+7l = 8x + 16 or | 4x +7| =4(8x+16) you can get an invalid x if the right side is negative with that x. (because the left side, with the abs value, can never be negative)

OpenStudy (phi):

to solve 1/4 (4x+7)= 8x+16 I would multiply both sides by 4 (who likes fractions!) 4x+7= 4*8x+4*16 collect terms and simplify

OpenStudy (anonymous):

Ok thank you so much!